# Physics-Guided Variational Model for Unsupervised Sound Source Tracking

### v1

@@ Abstract
Sound source tracking is often performed using classical array-processing algorithms. Alternative methods, such as machine learning, rely on ground truth position labels, which are costly to obtain. We propose a variational model that can perform single-source unsupervised sound source tracking in latent space, aided by a physics-based decoder. Our experiments demonstrate that the proposed method surpasses traditional baselines and achieves performance and computational complexity comparable to state-of-the-art supervised models. We also show that the method presents substantial robustness to altered microphone array geometries and corrupted microphone position metadata. Finally, the method is extended to multi-source sound tracking and the basic theoretical changes are proposed.
@@ 1 Introduction
Sound source localization or tracking consists of estimating the position or direction of arrival of one or more sources based on signals recorded by a multi‑microphone array. Traditionally, this is achieved using classical array signal processing methods, such as evaluating candidate directions [28, 7], analyzing system eigenstructures [27], or fitting parametric array models [29]. These methods, however, exhibit several limitations, including the need for grid searches, precise array calibration, and a strong dependence on initialization.
@@
With the growing interest in machine learning, neural network-based methods have been proposed as alternatives for the source localization and tracking problems [15, 4, 1]. These approaches use different architectures, depending on different input features and training objectives (classification or regression). Toward the state-of-the-art, we can find architectures that are more tailored to the problem of sound source tracking, either employing multi-dimensional convolutional layers [6] or parallelizing its encoder for more efficient and general operation [14]. The neural network-based approaches use supervised training, i.e., they depend on precise ground truth source position values. Obtaining such labels can be costly in both resources and time, also impeding training or fine-tuning to be done on-device.
@@
Alternatively, unsupervised learning–based approaches could be used to bridge this gap and eliminate the need for labeled data, while potentially offering more precise estimates than classical methods. However, most existing approaches are not fully unsupervised [16, 24] or rely on image or video modalities to facilitate tracking [32, 20]. The one existing fully unsupervised method [18] requires a very high computational cost and a large number of parameters, making it infeasible for most audio devices, and is further limited by its ad‑hoc design choices.
@@
In contrast, we propose a physics‑guided variational model for single‑source sound tracking. Our method is built upon a variational autoencoder structure, inheriting its underlying theoretical foundations. The physics‑based decoder injects spatial information into the latent space via backpropagation, enabling unsupervised estimation of source directions. We compare the variational model against classical and state‑of‑the‑art approaches across three experiments using real‑world data, achieving performance comparable to supervised learning‑based models. Furthermore, we demonstrate that our approach generalizes well to microphone array geometries that differ between training and testing. Compared to state‑of‑the‑art methods, it also exhibits substantial robustness to corrupted metadata. Finally, we derive a multi‑source extension of the proposed method and outline the theoretical framework required for its implementation. In this paper, however, we focus on the single‑source case, as the multi‑source implementation requires significant additional work and would shift the scope of the manuscript.
@@ 2 Sound source tracking
where ($\ast$) denotes the convolution between a room impulse response $\mathbf{h}_{nm}(t)$, from source $n$ to microphone $m$, and the signal produced by a source $n$ ($\mathbf{s}_{n}$). The number of sources $n$ is limited to $N$, and the sensor noise is denoted by $\mathbf{\epsilon}_{m}(t)$. The objective is to track the position of source $\mathbf{s}_{n}$, at discrete time $t$. Moreover, we assume that the sources are in the far-field of the array, i.e., the distance between microphones is significantly shorter than the distance between sources to the array [13]. The sound wave is thus treated as a plane wave without defined origin, and we only estimate its direction of arrival (DOA), described by an azimuth and an elevation angle. While this study restricts itself to far‑field conditions, the underlying methodology is compatible with near‑field extensions, which we leave for future work. Furthermore, we tackle the problem of tracking a single source ($N=1$), using unsupervised machine learning. Relevant existing methods related to the problem are explored next.
@@ 2.1 Classical
In this work, we use the steered response power (SRP) method [7, 13] as a baseline, as it is a representative spatial signal‑processing approach. SRP estimates the DOA by evaluating a predefined grid of candidate directions and assigning each direction a score that reflects how well the microphone signals align under that hypothesized geometry. It can be interpreted as a spatial projection of pairwise microphone cross‑correlations, using the time delays predicted by each candidate location. Common features used within SRP include raw cross‑correlation, generalized cross‑correlation (GCC), and its weighted variant, the GCC with phase transform (GCC‑PHAT). The latter normalizes the cross‑spectrum by the magnitude, emphasizing phase information, which is more robust at low signal‑to‑noise ratios (SNRs). On the negative side, the resolution of SRP depends on the granularity of the direction grid and, in contrast to probabilistic approaches such as maximum likelihood, it does not model uncertainty.
@@ 2.2 Supervised learning
Among supervised approaches, two methods can be distinguished as representing the state of the art: Cross3D [6] and Neural‑SRP [14]. The Cross3D method was proposed for single‑source DOA tracking, taking an SRP‑produced map as input. Its architecture is a 3‑dimensional CNN, where each dimension corresponds to elevation angle, azimuth angle, and time. The Neural‑SRP model, on the other hand, is able to generalize with respect to both the number and the positions of the microphones in the array. It processes pairwise GCC‑PHAT inputs using parallel copies of the same convolutional‑recurrent encoder. The encoder outputs are summed and passed through an MLP, which produces the estimated DOA. Both methods are trained with simulated data generated using the image source method [3] and tested on a real‑world dataset.
@@ 2.3 Unsupervised learning
We are, however, interested in fully unsupervised machine‑learning‑based methods that rely only on audio information for sound source localization or tracking. Such methods are, unfortunately, very scarce in the literature and exhibit strong limitations. To the best of our knowledge, the approach described next is the only prior work that is fully unsupervised. A time‑domain unsupervised learning‑based method is proposed in [18], using an autoencoder neural network. Multiple encoders recover a latent source signal, followed by paired decoders that model the forward acoustic transfer. The training loss is based on the reconstruction quality of the acoustic transfers and the cross‑channel consistency among encoder outputs. This method, however, results in a very high computational cost and parameter count due to its multiple encoders and decoders, driven by the need to reconstruct sound source signals. Next, we describe our proposed solution, which adopts a more feasible, lower‑complexity design.
@@ 5.3 Hyperparameters
5.4 Results
When the models are trained with auralized AWGN (Experiment 1), we observe that the proposed unsupervised approach outperforms the SRP-PHAT and is outperformed by the supervised-learning models of Cross3D and Neural-SRP, for both simulated and real-world data. This is an expected order of performance, since the classical approach is limited by not modeling uncertainty nor being trainable, and the methods with supervised learning had access to the ground truth values of direction of arrival.
@@ 5.4 Results
Conversely, we observe a different behavior in Experiment 3. By sampling the microphone positions from a Gaussian distribution around the LOCATA NAO robot’s array, we are able to evaluate the effect of uncertain microphone positioning on DOA estimation, i.e., training with metadata or data that differ from those used in evaluation. In this experiment, the proposed unsupervised approach is able to outperform the Cross3D method when evaluated on real‑world data, achieving performance comparable to Neural‑SRP. While both the proposed model and Neural‑SRP rely on GCC‑PHAT inputs and have access to metadata, the Cross3D method depends on an SRP map. Its reduction in performance likely occurs because it inherits limitations from the SRP method – such as grid dependence – which create a biased input that is further degraded by uncertain microphone positions. Cross3D also appears to have a lower capacity to generalize, given its substantially higher error on the LOCATA dataset compared to simulated data.
Furthermore, the real‑world performance of the proposed variational model improves relative to the other experiments. This counterintuitive behavior may arise from the additional excitation introduced by varying microphone positions – similar to system identification problems [21] – combined with the uncertainty modeling inherent to the probabilistic latent space (via reparameterization) and the decoder (through probabilistic modeling of time delays). For methods lacking explicit uncertainty modeling (Cross3D and Neural‑SRP), performance worsens for both simulated and real‑world data.
As a direct comparison between the proposed unsupervised approach and Neural‑SRP, we corrupt the metadata during evaluation at different levels. By applying AWGN to the metadata $(\mathbf{v}_{i},\mathbf{v}_{j})$ fed to the model, we cause it to differ from the real microphone positions. In practice, this is equivalent to a scenario where the system undergoes an initial calibration to define metadata values, but the microphone positions in the array are later changed without recalibration. We define the corruption percentage as the percentage of the maximum absolute value of $\mathbf{v}$ by which the unit‑variance AWGN is scaled. From Table 5, we see that as metadata corruption increases, the root‑mean‑square angular error of the proposed approach grows much less than that of the Neural‑SRP model, and the unsupervised approach outperforms Neural‑SRP from the 30% corruption level onward. A plausible explanation is that the proposed approach is based on variational autoencoders, inheriting their generalization properties through modeling the underlying data distribution within a known latent distribution.
@@
5.5 Complexity
Finally, we evaluate the complexity of the proposed model in terms of parameters, multiply-accumulate operations (MACs), and real time factor (RTF), which is shown in Table 6. Importantly, the decoder is not taken into account for the proposed model as it is unnecessary during inference mode. Also note that we obtain the RTF by dividing the calculation time for one step (hop length) by the hop length itself, averaged over 2000 seconds of audio. Since the proposed model has a similar architecture to that of the Neural-SRP, its number of parameters, MACs, and RTF are very similar to the supervised baseline. The Cross3D, on the other hand, presents much more parameters but less MACs and RTF, which happens because the model does not include any recurrent layer, mainly convolutional layers.
5.6 Multi-source scenario
$q_{\phi}(\mathbf{z}|\mathbf{g})=\sum_{n=1}^{N}\pi_{n}\,\mathrm{vMF}\!\left(\mathbf{z}|\boldsymbol{\mu}_{\phi,n},\kappa_{\phi,n}\right),$ (26)
@@ 6 CONCLUSIONS
The proposed method is evaluated over three experiments, which differ in noise type and uncertainty of testing geometry, compared to a classical and two supervised learning-based methods. The physics-guided variational model consistently outperforms the classical approach and presents comparable performance, in terms of root mean square angular error, to its supervised counterparts. Importantly, it retains higher generalization capacity given its variational nature, outperforming one of the supervised learning approaches when trained with array geometries that differ from evaluation. The proposed approach is also more robust to uncertainty in the microphone array metadata when compared to a state-of-the-art supervised model. Additionally, we analyze the models’ complexity showing that the proposed approach has similar complexity to that of the state-of-the-art.
Beyond this paper’s scope, we suggest the extension of the proposed model to a multi-source scenario. In Section 5.6, we derive the basic theory for the multi-source operation, which can be used as a starting point for future work. Furthermore, another extension would be the generalization to number of microphones, leveraging the parallel architecture as done in [14]. To maintain unsupervised operation, we suggest the proposed encoder to be extended as the combination of a continuous and categorical encoder, similar to [9]. The categorical part should treat the number of microphones as a random variable. The continuous encoder and physics-based decoder can inherit most of this paper’s proposed model, with minor changes.

### v2

@@ Abstract
Sound source tracking is commonly performed using classical array-processing algorithms, while machine-learning approaches typically rely on precise source position labels that are expensive or impractical to obtain. This paper introduces a physics-guided variational model capable of fully unsupervised single-source sound source tracking. The method combines a variational encoder with a physics-based decoder that injects geometric constraints into the latent space through analytically derived pairwise time-delay likelihoods. Without requiring ground-truth labels, the model learns to estimate source directions directly from microphone array signals. Experiments on real-world data demonstrate that the proposed approach outperforms traditional baselines and achieves accuracy and computational complexity comparable to state-of-the-art supervised models. We further show that the method generalizes well to mismatched array geometries and exhibits strong robustness to corrupted microphone position metadata. Finally, we outline a natural extension of the approach to multi-source tracking and present the theoretical modifications required to support it.
@@ 1 Introduction
Sound source localization or tracking consists of estimating the position or direction of arrival of one or more sources based on signals recorded by a multi-microphone array. Traditionally, this is achieved using classical array signal processing methods, such as evaluating candidate directions [29, 7], analyzing system eigenstructures [28], or fitting parametric array models [30]. These methods, however, exhibit several limitations, including the need for grid searches, precise array calibration, and a strong dependence on initialization.
@@
With the growing interest in machine learning, neural network-based methods have been proposed as alternatives for the source localization and tracking problems using, e.g., multilayer perceptron architectures [15], convolutional neural networks [4], and convolutional-recurrent networks [1]. The different structures also consider different input features and training objectives (classification or regression). Toward the state-of-the-art, we can find architectures that are more tailored to the problem of sound source tracking, either employing multi-dimensional convolutional layers [6] or parallelizing its encoder for more efficient and general operation [14]. The neural network-based approaches use supervised training, i.e., they depend on precise ground-truth source position values. Obtaining such labels can be costly in both resources and time, also impeding training or fine-tuning to be done on-device.
@@
Alternatively, unsupervised learning–based approaches could be used to bridge this gap and eliminate the need for labeled data, while potentially offering more precise estimates than classical methods. However, the existing approaches developed in this direction are not fully unsupervised, as they depend on clean speech data, forming a weakly supervised setup [16], or rely on a small set of labeled samples [25]. Some approaches that are unsupervised require image or video modalities to facilitate tracking [33, 21], which translates to additional sensors, like cameras [20], increasing the cost of audio/hearing devices. A different method that only depends on audio information [18] nevertheless needs simulated acoustic transfer functions during training, which require full knowledge of the array geometry and source positions. Moreover, the approach incurs very high computational cost and a large number of parameters, due to per-direction/per-microphone autoencoder models with four 1024-unit fully connected layers, making it infeasible for most audio devices, and it is further limited by ad-hoc design choices.
@@
In contrast, we propose a physics-guided variational model for single-source sound tracking. Our method is built upon a variational autoencoder structure, inheriting its underlying theoretical foundations. The physics-based decoder injects spatial information into the latent space via backpropagation, enabling unsupervised estimation of source directions. We compare the variational model against classical and state-of-the-art approaches across three experiments using real-world data, achieving performance comparable to supervised learning-based models. Furthermore, we demonstrate that our approach generalizes well to microphone array geometries that differ between training and testing. Compared to state-of-the-art methods, it also exhibits substantial robustness to corrupted metadata. Finally, we derive a multi-source extension of the proposed method and outline the theoretical framework required for its implementation. In this paper, however, we focus on the single-source case, as the multi-source implementation requires significant additional work and would shift the scope of the manuscript.
@@ 2 Sound source tracking
where ($\ast$) denotes the convolution between a room impulse response $\mathbf{h}_{nm}(t)$, from source $n$ to microphone $m$, and the signal produced by a source $n$ ($\mathbf{s}_{n}$). The number of sources $n$ is limited to $N$, and the sensor noise is denoted by $\mathbf{\epsilon}_{m}(t)$. The objective is to track the position of source $\mathbf{s}_{n}$, at discrete time $t$. Moreover, we assume that the sources are in the far-field of the array, i.e., the distance between microphones is significantly shorter than the distance between sources to the array [13]. The sound wave is thus treated as a plane wave without defined origin, and we only estimate its direction of arrival (DOA), described by an azimuth and an elevation angle. While this study restricts itself to far-field conditions, the underlying methodology is compatible with near-field extensions, which we leave for future work. Furthermore, we tackle the problem of tracking a single source ($N=1$), using unsupervised machine learning. Relevant existing methods related to the problem are explored next.
@@ 2.1 Classical
In this work, we use the steered response power (SRP) method [7, 13] as a baseline, as it is a representative spatial signal-processing approach. SRP estimates the DOA by evaluating a predefined grid of candidate directions and assigning each direction a score that reflects how well the microphone signals align under that hypothesized geometry. It can be interpreted as a spatial projection of pairwise microphone cross-correlations, using the time delays predicted by each candidate location. Common features used within SRP include raw cross-correlation, generalized cross-correlation (GCC), and its weighted variant, the GCC with phase transform (GCC-PHAT). The latter normalizes the cross-spectrum by the magnitude, emphasizing phase information, which is more robust at low signal-to-noise ratios (SNRs). On the negative side, the resolution of SRP depends on the granularity of the direction grid and, in contrast to probabilistic approaches such as maximum likelihood, it does not model uncertainty.
@@ 2.2 Supervised learning
Among supervised approaches, two methods can be distinguished as representing the state of the art: Cross3D [6] and Neural-SRP [14]. The Cross3D method was proposed for single-source DOA tracking, taking an SRP-produced map as input. Its architecture is a 3-dimensional CNN, where each dimension corresponds to elevation angle, azimuth angle, and time. The Neural-SRP model, on the other hand, is able to generalize with respect to both the number and the positions of the microphones in the array. It processes pairwise GCC-PHAT inputs using parallel copies of the same convolutional-recurrent encoder. The encoder outputs are summed and passed through an MLP, which produces the estimated DOA. Both methods are trained with simulated data generated using the image source method [3] and tested on a real-world dataset.
@@ 2.3 Unsupervised learning
We are, however, interested in fully unsupervised machine-learning-based methods that rely only on audio information for sound source localization or tracking. Such methods are, unfortunately, very scarce in the literature and exhibit strong limitations. To the best of our knowledge, the approach described next is the only prior work that is fully unsupervised. A time-domain unsupervised learning-based method is proposed in [18], using an autoencoder neural network. Multiple encoders recover a latent source signal, followed by paired decoders that model the forward acoustic transfer. The training loss is based on the reconstruction quality of the acoustic transfers and the cross-channel consistency among encoder outputs. This method, however, results in a very high computational cost and parameter count due to its multiple encoders and decoders, driven by the need to reconstruct sound source signals. Next, we describe our proposed solution, which adopts a more feasible, lower-complexity design.
@@ 5.3 Hyperparameters
5.4 Results
When the models are trained with auralized AWGN (Experiment 1), we observe that the proposed unsupervised approach outperforms the SRP-PHAT and is outperformed by the supervised-learning models of Cross3D and Neural-SRP, for both simulated and real-world data. This is an expected order of performance, since the classical approach is limited by not modeling uncertainty nor being trainable, and the methods with supervised learning had access to the ground-truth values of direction of arrival.
Conversely, we observe a different behavior in Experiment 3. By sampling the microphone positions from a Gaussian distribution around the LOCATA NAO robot’s array, we are able to evaluate the effect of uncertain microphone positioning on DOA estimation, i.e., training with metadata or data that differ from those used in evaluation. In this experiment, the proposed unsupervised approach is able to outperform the Cross3D method when evaluated on real-world data, achieving performance comparable to Neural-SRP. While both the proposed model and Neural-SRP rely on GCC-PHAT inputs and have access to metadata, the Cross3D method depends on an SRP map. Its reduction in performance likely occurs because it inherits limitations from the SRP method – such as grid dependence – which create a biased input that is further degraded by uncertain microphone positions. Cross3D also appears to have a lower capacity to generalize, given its substantially higher error on the LOCATA dataset compared to simulated data.
@@ 5.4 Results
Furthermore, the real-world performance of the proposed variational model improves relative to the other experiments. This counterintuitive behavior may arise from the additional excitation introduced by varying microphone positions – similar to system identification problems [22] – combined with the uncertainty modeling inherent to the probabilistic latent space (via reparameterization) and the decoder (through probabilistic modeling of time delays). For methods lacking explicit uncertainty modeling (Cross3D and Neural-SRP), performance worsens for both simulated and real-world data.
As a direct comparison between the proposed unsupervised approach and Neural-SRP, we corrupt the metadata during evaluation at different levels. By applying AWGN to the metadata $(\mathbf{v}_{i},\mathbf{v}_{j})$ fed to the model, we cause it to differ from the real microphone positions. In practice, this is equivalent to a scenario where the system undergoes an initial calibration to define metadata values, but the microphone positions in the array are later changed without recalibration. We define the corruption percentage as the percentage of the maximum absolute value of $\mathbf{v}$ by which the unit-variance AWGN is scaled. From Table 5, we see that as metadata corruption increases, the root-mean-square angular error of the proposed approach grows much less than that of the Neural-SRP model, and the unsupervised approach outperforms Neural-SRP from the 30% corruption level onward. A plausible explanation is that the proposed approach is based on variational autoencoders, inheriting their generalization properties through modeling the underlying data distribution within a known latent distribution.
@@
5.5 Complexity analysis
We evaluate the complexity of the proposed model in terms of parameters, multiply-accumulate operations (MACs), and real time factor (RTF), which is shown in Table 6. Importantly, the decoder is not taken into account for the proposed model as it is unnecessary during inference mode. Also note that we obtain the RTF by dividing the calculation time for one step (hop length) by the hop length itself, averaged over 2000 seconds of audio. Since the proposed model has a similar architecture to that of the Neural-SRP, its number of parameters, MACs, and RTF are very similar to the supervised baseline. The Cross3D, on the other hand, presents much more parameters but less MACs and RTF, which happens because the model does not include any recurrent layer, mainly convolutional layers.
5.6 Posterior uncertainty analysis
In Figure 4, we examine the behavior of the posterior concentration parameter under different reverberation and noise conditions. When the SNR is fixed and RT60 is varied (red curves), $\kappa_{\phi}$ shows a clear non-monotonic trend. For short RT60 values, the direct path dominates the GCC-PHAT structure, producing a higher peak in $p(\tau_{k}|\mathbf{g}_{k})$ and resulting in larger $\kappa_{\phi}$, i.e., higher confidence in the inferred direction. As RT60 increases to moderate values, early reflections introduce competing peaks around the direct-path delay, making the time-delay distribution less clearly uni-modal. Since the physics-guided decoder models each pairwise delay with a uni-modal likelihood, the encoder reflects this increased ambiguity by reducing the concentration value. At high RT60, the reverberant field becomes more diffuse and the structured secondary peaks weaken, causing the dominant mode to stabilize again. In this regime, $\kappa_{\phi}$ increases slightly, indicating that the model recovers part of its directional certainty despite the overall broader input distributions.
On the other hand, when RT60 is fixed and SNR is varied (black curves), the behavior of the concentration parameter becomes more monotonic. At high SNR, the inter-microphone phase structure is preserved, yielding sharp GCC-PHAT peaks and larger concentration values. As SNR decreases, additive noise progressively masks the phase information and flattens the delay distribution, reducing the contrast between the direct path and surrounding delays. The posterior then widens to remain consistent with the physics-guided likelihood, producing steadily smaller $\kappa_{\phi}$. Unlike the reverberation case, low SNR does not create structured competing peaks, simply decreasing signal coherence. As a result, the decline in concentration with SNR is smooth and directly reflects noise-induced uncertainty in the latent DOA estimate.
(a) Fixed RT60 of 0.1 s, SNR of 30 dB
(b) Fixed RT60 of 0.4 s, SNR of 15 dB
(c) Fixed RT60 of 0.8 s, SNR of 0 dB
Figure 4: Concentration $\kappa_{\phi}$ obtained with one model trained as in Experiment 1, with (upper axis, red) an SNR sweep at fixed RT60, and (lower axis, black) an RT60 sweep at fixed SNR, for different fixed conditions over the simulated test set.
6 Multi-source extension guidelines
$q_{\phi}(\mathbf{z}|\mathbf{g})=\sum_{n=1}^{N}\pi_{n}\,p\!\left(\mathbf{z}|\boldsymbol{\mu}_{\phi,n},\kappa_{\phi,n}\right),$ (26)
@@ 7 Conclusions
The proposed method is evaluated over three experiments, which differ in noise type and uncertainty of testing geometry, compared to a classical and two supervised learning-based methods. The physics-guided variational model consistently outperforms the classical approach and presents comparable performance, in terms of root mean square angular error, to its supervised counterparts. Importantly, it retains higher generalization capacity given its variational nature, outperforming one of the supervised learning approaches when trained with array geometries that differ from evaluation. The proposed approach is also more robust to uncertainty in the microphone array metadata when compared to a state-of-the-art supervised model. Additionally, we analyze the models’ complexity – showing that the proposed approach has similar complexity to that of the state-of-the-art – and the behavior of the estimated concentration parameter, connecting its interpretation to the notion of uncertainty.
Beyond this paper’s scope, we suggest the extension of the proposed model to a multi-source scenario. In Section 6, we derive the basic theory for the multi-source operation, which can be used as a starting point for future work. Furthermore, another extension would be the generalization to number of microphones, leveraging the parallel architecture as done in [14]. To maintain unsupervised operation, we suggest the proposed encoder to be extended as the combination of a continuous and categorical encoder, similar to [9]. The categorical part should treat the number of microphones as a random variable. The continuous encoder and physics-based decoder can inherit most of this paper’s proposed model, with minor changes.
@@ Acknowledgment
This work was supported by the Robust AI for SafE (radar) signal processing (RAISE) collaboration framework between Eindhoven University of Technology and NXP Semiconductors, including a Privaat-Publieke Samenwerkingen-toeslag (PPS) supplement from the Dutch Ministry of Economic Affairs and Climate Policy.